OpenStudy (anonymous):

Theorem: For any given area A, the rectangle that has the least perimeter is a square. Part I: Complete the proof that is outlined below. Submit each step as part of your final answer. Proof: In terms of the side length, x, the formula for the perimeter of a square: P= 2x+((2A)/x) Define a new variable: v=sqrt(x)-(sqrt(A))/(sqrt(x)) In terms of the new variable, v, compute and simplify: 2v^2 = _____. Rewrite this equation to get the formula for P alone on one side. Replace the formula by the variable P. Write that result by completing the equation below: P = 2v^2

OpenStudy (anonymous):

Let the given area be A, this is constant length and breadth of rectangle be given by l and b respectively then the Perimeter is\[P=2(l+b)\] This equation contains 2 variables, we shall eliminate one using the constant we are given \[A=lb\]\[b=\frac{A}{l}\]Then P is a function of l is\[P(l)=2(l+\frac{A}{l})\] Use the second derivative test, find \[\frac{dP}{dl}\] equate it to 0, find the value of l for which there is a critical point, then find \[\frac{d^2P}{dl^2}\] then evaluate the second derivative for that value of l which u calculated in first derivative, if the value is less than 0, perimeter is maximum, otherwise it is minimum

OpenStudy (anonymous):

using the results from 2nd order test ull be able to prove what is required

OpenStudy (anonymous):

@Nishant_Garg I love you. You are my favorite person in the entire world right now haha. Thank you so much!!!

OpenStudy (anonymous):

Haha, thank me after you've solved the question rather! :)